# Enabling AI Scientists to Recognize Innovation: A Domain-Agnostic Algorithm for Assessing Novelty

### v2

@@ 5.1 Why a Non-LLM Novelty Assessment Algorithm is Necessary?
Assessing the novelty of a research idea is inherently difficult, subjective, and resource-intensive. While LLMs have the potential to assist in this process, their effectiveness is limited by the challenges outlined in the Introduction. Our experiments (see Table 2) highlight these issues: without an integrated search tool, even the most advanced reasoning models’ performance was comparable to random guessing (AUROC=̃0.5). When a search tool was introduced, Sonnet-3.7 achieved similar accuracy on the NeurIPS test set (AUROC=̃0.8) but experienced significant degradation (AUROC=̃0.6) on both the Nature Medicine and cross-domain test sets.
In contrast, our proposed RND algorithm can produce more reliable and consistent results, as seen in Table 2. Our algorithm is better at distinguishing genuinely novel ideas from the large pool of candidates from mixing research domains (AUROC=̃0.78 v.s Other’s AUROC<=0.6). Such cross-domain novelty assessment capability is crucial to AI scientist, as more and more innovation happened in inter-discipline of research domains.
Recent advancements in reinforcement learning for reasoning models, such as those demonstrated in Deepseek-R1 (Guo et al. (2025)), suggest the potential of rule-based reward systems to guide model development. Our RND algorithm could serve as a sophisticated rewarding mechanism, potentially enhancing reasoning model’s capabilities in generating novel scientific ideas and advancing the role of AI in scientific innovation.
5.2 Accuracy in Each Domain
In NeurIPS test set, the neighbor density-based RND algorithm, absolute local density-based HD algorithm and Sonnet-3.7 with literature search tools achieved AUROC better than 0.8. When it comes to another domain (Nature Medicine test set for biomedical research), only RND and HD achieved AUROC at approximately 0.7; the other algorithms, including reasoning model such as Sonnet-3.7, degraded to AUROC 0.6. The strong performance of HD in the two respective domains suggests that measuring the local density (average semantic distance between a given idea and its nearest neighbors) in the historical literature database can effectively indicate the novelty of that idea in a single research domain. Since our proposed algorithm also incorporates semantic density, it exhibited similar accuracy level.
@@ 5.3 Domain-invariant Accuracy
However, when tested in the cross-domain test set (the Mixed test set), which includes ideas from both computer science and biomedicine, the performance of HD significantly degraded, with its AUROC dropping to 0.362. In contrast, the AUROC of our proposed algorithm remained robust at 0.795, similar to its performance in the single-domain test sets. As demonstrated in the Table 3, the relative position of idea’s local density among all of its neighbor’s local density is crucial for comparing novelty across different domains.
We argue that RND algorithm hold domain-invariant property, i.e. the distribution of novelty scores produced by RND is identical regardless of the tested domain, which explained why our relative density-based approach succeeds in cross-domain scenarios. According to the mathematical reasoning in Appendix A.3, we concluded the distribution of novelty score $S$ is only subject to $P$ (the number of neighbors considered); thus it is invariant to the validation domain. Furthermore, in Figure 1 (right panel) and 5, the actual distribution of scores echoed the theoretical analysis. Such domain-invariant property is crucial for conducting multi-disciplinary scientific research, where ideas from diverse fields must be compared and evaluated effectively.
5.4 Why Validation Methods Differ Between Novel and Non-Novel?
When building our test set, an obvious approach might be to use symmetrical sources - for example, using accepted NeurIPS papers as novel samples and rejected NeurIPS papers (specifically those rejected for lacking novelty) as non-novel samples. However, this approach presents significant limitations. Firstly, very few top-tier venues publicly release review comments with explicit novelty assessments, making such data scarce and difficult to generalize across domains. Secondly, papers may be rejected for "lack of novelty" due to incremental advances or methodological similarities, even when addressing previously unexplored topics.
Instead, our definition of novelty relies on how extensively similar ideas have been studied in the literature. Following this definition, we selected highly-cited papers from recent years as our non-novel samples, as these papers represent ideas that have been thoroughly explored and extended by numerous subsequent works. While high citation count itself can indicate either novelty or utility, papers that are both recent and highly-cited typically represent research areas that have quickly become crowded with similar work, making the original contributions less novel by our working definition. Further details on our sampling methodology can be found in Section 4.1.
@@ 5.5 Limitations & Future Work
Several limitations of our work warrant further exploration.
First, the algorithm relies heavily on large-scale literature databases with semantic embeddings. Biases in the literature database could potentially influence novelty assessments, especially if certain areas of research are underrepresented or if publication biases exist within fields.
Second, the algorithm’s performance is also dependent on the quality of semantic embeddings for representing complex scientific concepts. While the M3 model demonstrated effectiveness, domain-specific fine-tuning could potentially improve performance. Future work should investigate specialized embedding models for scientific literature that better capture the complex semantics of scientific abstracts, particularly for technical terminology and methodological nuances.
Third, our validation methodology, while avoiding the need for expert labeling, relied on non-novel samples that may be too easily distinguishable from novel ones. By using historical highly-cited papers as non-novel examples, rather than borderline cases such as recently rejected papers or incremental work from current journals, we created a simplified assessment scenario compared to the subtle distinctions scientists face in real research settings. However, the fact that none of the tested algorithms achieved saturated AUROCs even in this relatively straightforward scenario demonstrates the fundamental challenge of novelty assessment and validates our comparative analysis.
Looking ahead, we envision several promising directions for future work:
@@
2. Enhancing Reasoning Model: As highlighted in our discussion, current reasoning models struggle with reliable novelty assessment across domains. We propose utilizing our RND algorithm as a sophisticated reward mechanism within reinforcement learning frameworks for training reasoning models for AI research. By providing domain-invariant novelty signals during training, we could potentially guide models to generate more innovative scientific ideas while maintaining scientific validity.
These advancements would further enhance AI’s role in scientific research by accelerating idea generation, refining research hypotheses, and potentially uncovering interdisciplinary connections that might otherwise remain unexplored.
@@
Appendix A Algorithm Analysis
A.1 Effects of Parameter P
@@ 3. Variance and Sensitivity with Finite $P$
  • Smaller $\textbf{Smaller }P:$ The variance $\operatorname{Var}(F_{P}(ND_{i}))$ is larger, leading to a noisier (less reliable) estimation of the quantile, and hence of the novelty score.
  • Larger $\textbf{Larger }P:$ The variance decreases, yielding a more accurate estimation of the true quantile $F(ND_{i})$.
@@ A.2 Effects of Parameter Q
A.2.1 Interpretation of Variance Scaling
@@
A.3 Domain-Invariant
A.3.1 Theoretical Analysis
@@
Thus, when $P$ changes (e.g., $P=50$, $100$, or $500$), the change in variance—and hence the fluctuation in the score—is proportional to $\frac{1}{P}$ and is independent of the corpus. In other words,

### v3

@@ 5.1 Why a Non-LLM Novelty Assessment Algorithm is Necessary?
While LLMs have the potential to assess novelty, the reliability of their judgment is limited, as outlined in the Introduction section (1). Our experiments (see Table 2) echoed such concern: without an integrated search tool, even the most advanced reasoning models’ performance was comparable to random guessing (AUROC=̃0.5). When a search tool was introduced, Sonnet-3.7 achieved similar accuracy on the NeurIPS test set (AUROC=̃0.8) but experienced significant degradation (AUROC=̃0.6) on both the Nature Medicine and cross-domain test sets.
In contrast, our proposed RND algorithm can produce more reliable and consistent results, as seen in Table 2. Additionally, our algorithm is better at distinguishing genuinely novel ideas from the large pool of candidates from mixing research domains (AUROC=̃0.78 v.s Other’s AUROC<=0.6). Such cross-domain novelty assessment capability is crucial to AI scientist, as more and more innovation happens in the intersection of research domains.
@@ 5.2 Domain-invariant Accuracy
The most significant advantage of RND algorithm is in being able to accurately compare novelty across different domains.
As demonstrated in the Table 3, though local density-based HD algorithm and LLM with literature search tools achieved 0.8+ AUROC in NeruIPS test set, their performance degraded in Nature Medicine test set (0.699 for HD, 0.673 for Deepseek-r1). Moreover, when tested in the cross-domain test set (the Mixed test set), the AUROC of HD and LLM significantly degraded below 0.6. In contrast, the AUROC of our proposed RND algorithm remained robust (0.820, 0.765, 0.795) across each single domain and mixed domain test sets.
We argue that such benefits come from RND algorithm’s domain-invariant property, i.e. the distribution of novelty scores produced by RND is identical regardless of the tested domain, which explained why our relative density-based approach succeeds in cross-domain scenarios. According to the mathematical reasoning in Appendix A.4, we concluded the distribution of novelty score $S$ is only subject to $P$ (the number of neighbors considered); thus it is invariant to the validation domain. Furthermore, in Figure 2 (right panel) and 6, the actual distribution of scores echoed the theoretical analysis. Such domain-invariant property is crucial for conducting multi-disciplinary scientific research, where ideas from diverse fields must be compared and evaluated effectively.
@@ 5.3 Limitations & Future Work
First, the RND algorithm’s effectiveness is inherently linked to the breadth and quality of the underlying literature database and semantic embedding model. Incomplete literature coverage or suboptimal embedding quality can impact novelty assessment accuracy. While implementing RND requires more infrastructure than LLM-as-judge approaches, this investment enables consistent, deterministic assessments that remain stable across model versions and prompting strategies.
Second, in the test set created by our validation methodology, while the need for expert labeling was avoided, the non-novel samples may be too easily distinguishable from novel ones. By using historical highly-cited papers as non-novel examples, rather than borderline cases such as recently rejected papers or incremental work from current journals, we created a simpler assessment scenario than the actual scenario in real research settings. However, the fact that none of the tested algorithms achieved saturated AUROCs even in this relatively straightforward scenario demonstrates the fundamental challenge of novelty assessment.
While acknowledging these limitations, we believe our work establishes a solid foundation for several promising research directions that could significantly advance automated scientific discovery.
@@
2. Enhancing Reasoning Capability for LLM: We propose utilizing our RND algorithm as a sophisticated reward mechanism within reinforcement learning frameworks for training reasoning models for AI research. By providing domain-invariant novelty signals during training, it could potentially guide models to generate more innovative scientific ideas while maintaining scientific validity.
@@
Appendix A Algorithms
A.1 Pseudocode
Algorithm 1 Find Neighbors and Calculate Neighbor Density
1:function Neighbor(Input, P, Q)
3:  C $\leftarrow$ []
5:  $neighbors\_for\_count\leftarrow neighbors[:Q]$ $\triangleright$ Only use the Q nearest neighbors to calculate density
6:  $neighbors\_for\_distribution\leftarrow neighbors[:P]$ $\triangleright$ The P nearest neighbors are used to calculate distribution
7:  for each paper in neighbors_for_countneighbors\_for\_count do
8:   $\mathbf{v}_{paper}\leftarrow$ Get_Embedding(paper)
10:   C.Append(distance)
11:  end for
12:  $ND_{Input}\leftarrow$ mean(C)
13:  return $ND_{Input}$, neighbors_for_distributionneighbors\_for\_distribution
14:end function
Algorithm 2 Calculate Novelty Score of Given Idea
1:Input: Idea
2:Output: A score in the range of 0 to 100
3:D $\leftarrow$ []
[…]
5:for paper in neighbors do
6:  $ND_{paper}$, _ $\leftarrow$ Neighbor(paper, P, Q)
7:  D.Append($ND_{paper}$)
8:end for
9:score $\leftarrow\frac{|\{ND\in D\mid ND\leq ND_{Idea}\}|}{|D|}\times 100$
10:Return score
A.2 Effects of Parameter P
@@ 3. Variance and Sensitivity with Finite $P$
  • $\textbf{Smaller }P:$ The variance $\operatorname{Var}(F_{P}(ND_{i}))$ is larger, leading to a noisier (less reliable) estimation of the quantile, and hence of the novelty score.
  • $\textbf{Larger }P:$ The variance decreases, yielding a more accurate estimation of the true quantile $F(ND_{i})$.
@@ A.3 Effects of Parameter Q
A.3.1 Interpretation of Variance Scaling
@@
A.4 Domain-Invariant
A.4.1 Theoretical Analysis
@@
Thus, when $P$ changes (e.g., $P=50$, 100100, or $500$), the change in variance—and hence the fluctuation in the score—is proportional to $\frac{1}{P}$ and is independent of the corpus. In other words,
@@ D.3 Comparison of AUROC of RND algorithm with different Q values
Appendix E Computing Resources
To conduct our experiment, the necessary computing resources includes
1. Literature database with embedding vector: we used a server with 16 CPU cores, 128GB RAM, 2TB SSD disk(which actually used around 1TB) to deploy an ElasticSearch version 8 as literature search engine. The embedding of each literature is obtained using public available API services.
2. Algorithm development and experiment: a server with 16 CPU cores, 128GB RAM.